Benvenuto is a pizzeria operating in a competitive market. Its short-run production function is shown below. Benvenutos weekQuestion 6 Homework . Unanswered In the scenario above, suppose Benvenuto currently produces 438 pizzas per week. What is theQuestion 7 Homework • Unanswered In the scenario above, when Benvenuto produces 924 pizzas, what is its average fixed cost ($Question 8 Homework • Unanswered In the scenario above, when Benvenuto produces 1,075 pizzas, what is its average variable coQuestion 9 Homework • Unanswered In the scenario above, when Benvenuto produces 1,168 pizzas, what is its average total cost
Answer #1

FC=700

VC=450*L+6*Q

TC=FC+VC

AFC=FC/Q

AVC=VC/Q

ATC=TC/Q

MC=change in T/change in Q

L    Q    FC    VC    TC    AFC    AVC ATC    MC
1 52 700 762 1462 13.46 14.65 28.12 14.65
2 172 700 1932 2632 4.07 11.23 15.30 9.75
3 341 700 3396 4096 2.05 9.96 12.01 8.66
4 536 700 5016 5716 1.31 9.36 10.66 8.31
5 738 700 6678 7378 0.95 9.05 10.00 8.23
6 924 700 8244 8944 0.76 8.92 9.68 8.42
7 1075 700 9600 10300 0.65 8.93 9.58 8.98
8 1168 700 10608 11308 0.60 9.08 9.68 10.84
9 1184 700 11154 11854 0.59 9.42 10.01 34.13

6) TC for 438 units = 700+[4*450 + 6*438] = 700+4428 = 5128

Tc for 439 units = 700+[4*450 + 6*439] = 700+4434 = 5134

MC = 5134-5128 = 6

7) AFC = 0.76

8) AVC = 8.93

9) ATC = 9.68